# Multi-Token Attention

### v1

@@ Abstract
Soft attention is a critical mechanism powering LLMs to locate relevant parts within a given context. However, individual attention weights are determined by the similarity of only a single query and key token vector. This “single token attention” bottlenecks the amount of information used in distinguishing a relevant part from the rest of the context. To address this issue, we propose a new attention method, Multi-Token Attention (MTA), which allows LLMs to condition their attention weights on multiple query and key vectors simultaneously. This is achieved by applying convolution operations over queries, keys and heads, allowing nearby queries and keys to affect each other’s attention weights for more precise attention. As a result, our method can locate relevant context using richer, more nuanced information that can exceed a single vector’s capacity. Through extensive evaluations, we demonstrate that MTA achieves enhanced performance on a range of popular benchmarks. Notably, it outperforms Transformer baseline models on standard language modeling tasks, and on tasks that require searching for information within long contexts, where our method’s ability to leverage richer information proves particularly beneficial.
@@ 1 Introduction
We first experiment with a motivating toy task that reveals the shortcoming of standard attention and demonstrate that MTA can easily solve it. Next, we test our method at scale by pre-training 880M parameters models on 105B tokens on a standard language modeling task. There we see MTA bring improvements in terms in validation perplexity as well as standard benchmark tasks, while only increasing the number of parameters by $0.001\%$. Further, we evaluate the resulting models on long-context tasks such as Needle-in-the-Haystack and BabiLong where MTA outperforms the baselines by a significant margin.
@@ 2 Background on multi-head attention
Figure 1: Multi-Token Attention (MTA) on the right, compared to standard attention on the left. In MTA, within each head we apply a key-query convolution on the attention scores and, after softmax normalization, head convolution across groups of heads. Additionally, we apply a group normalization with depth-dependent scaling before final concatenation.
@@ 3 Multi-Token Attention
As shown in Figure 1 (right), our proposed Multi-Token Attention consists of three important components built on top of multi-head attention: key-query convolution, head mixing convolution, and group normalization with depth scaling. We propose key-query convolution to combine multiple keys and queries within heads, and head convolution to share knowledge between heads and amplify important information. Finally, we apply group normalization with depth scaling (Ye et al., 2024) to push back against residual streams and improve gradient flow. In this section, we will describe how MTA works.
@@ 3.3 Putting everything together
Finally, we follow Ye et al. (2024) and apply normalization with layer-dependent scaling for each head independently to improve gradient flow. We also report results without this step.
@@ 4 Experiments
We conduct experiments with the MTA architecture, comparing to several baselines on a set of standard and long-range dependency tasks, starting with a toy task. Unless otherwise specified, we apply key-query convolution pre-softmax and head mixing post-softmax.
@@ 4.1 Motivating toy task
The results are shown in Table 1. As expected, Transformer with standard multi-head attention struggles to solve this task, often completely failing to find the target blocks even though the questions have only $L=2$ letters. This highlights the inherent limitation of standard attention conditioned on single token vectors. For MTA, we set the query kernel size $c_{q}=2$ to match the number of question letters, and the key kernel $c_{k}=2N-1$ so it can cover a whole block on both sides. No head convolution is used. This way, it is sufficient for each query vector to encode only a single question letter while the convolution operation can aggregate their locations to find the target block. As a result, MTA successfully solves all the versions of the task with near zero error rate. See Appendix C for more training details.
@@ 4.2 Large language modeling
For language modeling experiments, we perform pre-training of 880M-size models and compare the Transformer model (Vaswani et al., 2017), Differential Transformer (Diff Transformer) (Ye et al., 2024), and Transformer with MTA. Diff Transformer calculates attention scores as the difference between two separate softmax attention maps, thus being related to our head convolution approach, so we use it as a baseline. All models are trained in the same setup on the SlimPajama (Soboleva et al., 2023) dataset for 105B tokens using the Lingua framework (Videau et al., 2024). Training details are provided in Appendix D. To improve training efficiency, we apply the key-query convolution on every 4th layer, while head convolution is applied on all layers. Kernel dimensions are fixed at $c_{q}=6,c_{k}=11$, and we mix groups of $c_{h}=2$ heads within each layer. We ablate these parameters in Section 4.6.
For each model, we conduct training twice and report average validation perplexity in Table 2. We observe consistent improvements across all validation datasets for the model trained with MTA, even though key-query convolution was only applied to a quarter of the layers, operating with less learnable parameters than Diff Transformer (see Appendix Table 9). We also note that group normalization with layer scaling is an important component that drives superior performance for both the Diff Transformer and MTA architectures.
@@
Diff w/o group norm. 4.63 13.41 20.06 14.34 4.26 10.06 11.59 11.19 (0.00)
Diff transformer 4.62 13.33 19.99 14.28 4.25 10.04 11.54 11.14 (0.02)
MTA w/o group norm. 4.63 13.36 20.03 14.31 4.25 10.05 11.52 11.16 (0.01)
MTA 4.60 13.26 19.90 14.22 4.22 09.96 11.44 11.09 (0.00)
@@ 4.4 Long-range dependency tasks
Needle-In-A-Haystack We probe our models by inserting 2, 4, and 6 needles in 2k and 4k context windows at varying depths. We additionally ensure that for each new sample, the needles are shuffled, thus removing bias the models might have toward extracting the needle that was inserted first or last. Each setup is evaluated on 500 samples. Accuracy evaluations are averaged across the depths of insertion. As reported in Table 5, we observe a significant improvement in needle extraction abilities for models trained with MTA across all needle counts and varying context lengths. Breakdown by depth is reported in Appendix G.
Pretraining, 2k cont. Long context finetuning, 4k cont.
Model N=2 N=4 N=6 N=2 N=4 N=6
Transformer 82.1 56.4 44.7 90.5 41.1 31.9
Diff transformer 96.0 62.5 55.6 95.7 73.8 60.0
MTA w/o group norm. 95.3 92.6 77.4 94.4 96.3 79.0
MTA 92.8 65.1 63.0 97.6 77.7 67.0
Table 5: Multi-needle retrieval accuracy (%) when varying the number of needles (N) averaged over different needle insertion depths. Pretrained models are evaluated with 2K context, while finetuned models are evaluated with 4K context.
@@ 4.5 Kernel patterns
Other kernels, for example, can be viewed as priming, amplifying if the same key was attended by previous queries (head 5 on Figure 8), or edge detecting, amplifying the first or last of multiple contiguous keys with high attention (heads 8 and 11 on Figure 10). We leave further exploration of key-query kernel maps and their meaning for future explorations.
@@ 4.6 Ablations
Key-query conv Head conv Group norm PPL $\downarrow$
$\checkmark$ $\checkmark$ $\checkmark$ 11.09
$\checkmark$ $\checkmark$ $\times$ 11.16
$\checkmark$ $\checkmark$ no scaling 11.13
$\checkmark$ $\checkmark$ layer-norm scaling 11.41
$c_{q}=4$, $c_{k}=9$$\times$ $\checkmark$ 11.23
$c_{q}=6$, $c_{k}=11$$\times$ $\checkmark$ 11.23
$c_{q}=8$, $c_{k}=13$$\times$ $\checkmark$ 11.31
post-softmax post-softmax $\checkmark$ 11.11
post-softmax post-softmax $\times$ 11.19
pre-softmax pre-softmax $\checkmark$ 11.10
pre-softmax pre-softmax $\times$ 11.20
@@ 5 Related work
There has been a number of attempts to focus the soft attention mechanism on important context. For example, modifications have been proposed to make softmax sharper. Martins & Astudillo (2016) propose to replace softmax with sparsemax for sparse activations. Veličković et al. (2024); Nakanishi (2025) propose Adaptive temperature and Scalable-Softmax that adjust exponential base in softmax attention. Irrelevant tokens can be altogether removed from memory (Sukhbaatar et al., 2021) to make attention easier. In Golovneva et al. (2024); Desrochers et al. (2024) the authors incorporate contextual information into position encodings to allow positions to be amplified by context. More recently, several noise canceling mechanisms have been proposed for attention. Diff transformer (Ye et al., 2024) uses differential amplifiers to focus attention to the relevant context, which is related to our head mixing step. Cang et al. (2025) further extends it by introducing normalization steps. OpAmp adaptation (Wu et al., 2025) propose adapters, which are fine-tuned with noisy context to enhance an LLMs’ denoising capability. Xu et al. (2024) modified attention so that keys and values can shift by one time step. This can be viewed as a special case of MTA where the convolution performs shifting in the key dimension.
@@
Convolution layers have been used with attention (Gehring et al., 2017; Wu et al., 2019), especially in vision Transformers, to decompose each image into a sequence of tokens with fixed length, and then apply multiple standard Transformer layers (Xiao et al., 2021; Wu et al., 2021). There have been some attempts to include convolution in language modeling as well. For example, Liu et al. (2018) use convolution to compress the keys and values in the multi-headed attention by a factor of 3 thus allowing to process sequences 3x in length. Liu & Lapata (2019) later augment this architecture with the ability to encode multiple documents in a hierarchical manner, and further improve on long-context summarization tasks. Later Subramanian et al. (2020) propose three hierarchical models to learn different levels of abstraction in language by interchanging casual transformer layers with convolutions or pooling, or compressing representations for tokens further away in the past. Gulati et al. (2020) proposed a Conformer architecture for speech recognition, where a convolution module is applied to multi-head self-attention output. This architecture was later used to encode speech in the Llama-3 herd of models (Grattafiori et al., 2024). However, none of these methods apply convolution to the attention weights.
@@ Appendix C Toy task details
Figure 5: Accuracy (%) on QA1-5 tasks in BabiLong benchmark. Note the random performance on QA3 is 16.67%, thus all models perform poorly on QA3. On other tasks, MTA demonstrate strong performance compared to baselines especially when there is lengthy distraction text (4K).
@@ Appendix D Language model training details
Training setup follows LLaMa architecture (Touvron et al., 2023), and includes RMSNorm pre-normalization (Zhang & Sennrich, 2019), SwiGLU activation function (Shazeer, 2020), and Rotary Embeddings (Su et al., 2024). Table 8 shows a detailed hyperparameter breakdown for 880M model pretraining. All models were trained in the same setup, except Diff Transformer, which had twice less heads to account for doubled dimension.
@@
Parameter Value
Dimension, $D$ $1536$
Layers $24$
Heads, $M$ $16$
RoPE theta $100,000$
Batch size (tokens) $262,144$
Context length $2048$
Learning rate $41.5e-$ 4
Weight decay $0.05$
Warm up steps $375$
Table 8: Hyperparameters for 880M Transformer training.

### v2

@@ Abstract
Soft attention is a critical mechanism powering LLMs to locate relevant parts within a given context. However, individual attention weights are determined by the similarity of only a single query and key token vector. This “single token attention” bottlenecks the amount of information used in distinguishing a relevant part from the rest of the context. To address this issue, we propose a new attention method, Multi-Token Attention (MTA), which allows LLMs to condition their attention weights on multiple query and key vectors simultaneously. This is achieved by applying convolution operations over queries, keys and heads, allowing nearby queries and keys to affect each other’s attention weights for more precise attention. As a result, our method can locate relevant context using richer, more nuanced information that can exceed a single vector’s capacity. Through extensive evaluations, we demonstrate that MTA achieves enhanced performance on a range of popular benchmarks. Notably, it outperforms Transformer baseline models on standard language modeling tasks, and on tasks that require searching for information within long contexts, where our method’s ability to leverage richer information proves particularly beneficial [^1] Code is available at https://github.com/facebookresearch/RAM/tree/main/projects/mta.
@@ 1 Introduction
We first experiment with a motivating toy task that reveals the shortcoming of standard attention and demonstrate that MTA can easily solve it. Next, we test our method at scale by pre-training 880M parameters models on 105B tokens on a standard language modeling task. There we see MTA bring improvements in terms in validation perplexity as well as standard benchmark tasks, while only increasing the number of parameters by $0.001\%$. Further, we evaluate the resulting models on long-context tasks such as Needle-in-the-Haystack and BabiLong where MTA outperforms the baselines by a significant margin. Finally, we investigate scaling laws of the proposed model and compare them with baselines.
@@ 2 Background on multi-head attention
Figure 1: Multi-Token Attention (MTA) (right), compared to standard attention (left). In MTA, within each head we apply a key-query convolution on the attention scores and a head convolution across groups of heads. We repeat this operation after the softmax. Finally, we apply a group normalization with scalar gating before final concatenation.
@@ 3 Multi-Token Attention
As shown in Figure 1 (right), our proposed Multi-Token Attention consists of three important components built on top of multi-head attention: key-query convolution, head mixing convolution, and group normalization with gating mechanism. The overall MTA convolution applies the key-query convolution to combine multiple keys and queries within heads, and the head convolution to share knowledge between heads and amplify important information. Finally, we apply group normalization with scalar gating to push back against residual streams and improve gradient flow. In this section, we will describe each component of MTA in detail.
@@ 3.3 Putting everything together
Finally, we follow Ye et al. (2024) who also apply normalization, but in our case we replace the layer-dependent scaling with a sigmoid gating mechanism. Sigmoid gating allows the model to switch heads on and off across layers and better adapt to different tasks. We also perform ablations across different normalization approaches.
@@ 4 Experiments
We conduct experiments with the MTA architecture, comparing to several baselines on a set of standard and long-range dependency tasks, starting with a toy task. Unless otherwise specified, we apply the MTA convolution both pre-softmax and post-softmax.
@@ 4.1 Motivating toy task
For this task, we train a small Decoder-only Transformer model with 4 layers, 2 heads and 256 hidden dimensions. The results are shown in Table 1. As expected, Transformer with standard multi-head attention struggles to solve this task, often completely failing to find the target blocks even though the questions have only $L=2$ letters. This highlights the inherent limitation of standard attention conditioned on single token vectors. For MTA, we set the query kernel size $c_{q}=2$ to match the number of question letters, and the key kernel $c_{k}=2N-1$ so it can cover a whole block on both sides. To simplify the experiments, the key-query convolution is only applied before softmax, and no head convolution is used. This way, it is sufficient for each query vector to encode only a single question letter while the convolution operation can aggregate their locations to find the target block. As a result, MTA successfully solves all the versions of the task with near zero error rate. See Appendix C for more training details.
@@ 4.2 Large language modeling
For language modeling experiments, we perform pre-training of 880M-size models and compare the Transformer model (Vaswani et al., 2017), Differential Transformer (Diff Transformer) (Ye et al., 2024), Transformer with Talking heads attention (Shazeer et al., 2020), and Transformer with MTA. Diff Transformer calculates attention scores as the difference between two separate softmax attention maps, while Talking heads attention applies linear projections across heads before and after softmax operations, thus being related to our head convolution approach, so we use them as baselines. All models are trained in the same setup on the SlimPajama (Soboleva et al., 2023) dataset for 105B tokens using the Lingua framework (Videau et al., 2024). Training details are provided in Appendix D. To improve training efficiency, we apply the key-query convolution on every 4th layer, while head convolution is applied on all layers. Kernel dimensions are fixed at $c_{q}=6,c_{k}=11$, and we mix groups of $c_{h}=16$ heads within each layer. We ablate these parameters in Section 4.6.
For each model, we conduct training twice and report average validation perplexity in Table 2. We observe consistent improvements across all validation datasets for the model trained with MTA, even though the key-query convolution was only applied to a quarter of the layers, keeping the total number of parameters on-par with the Transformer baseline (see Appendix Table 8). We also note that group normalization with layer scaling is an important component that drives superior performance for both the Diff Transformer and MTA architectures.
@@
Diff transformer 4.62 13.33 19.99 14.28 4.25 10.04 11.54 11.15 (0.02)
Talking heads 4.59 13.26 19.82 14.15 4.20 9.93 11.33 11.04 (0.00)
MTA 4.54 13.09 19.63 14.00 4.12 9.76 11.18 10.91 (0.01)
@@ 4.4 Long-range dependency tasks
Needle-In-A-Haystack We probe our models by inserting 2, 4, and 6 needles in 2k and 4k context windows at varying depths. We additionally ensure that for each new sample, the needles are shuffled, thus removing bias the models might have toward extracting the needle that was inserted first or last. Each setup is evaluated on 500 samples. Accuracy evaluations are averaged across the depths of insertion. As reported in Figure 3, we observe a significant improvement in needle extraction abilities for models trained with MTA across all needle counts and varying context lengths. Breakdown by depth is reported in Appendix G.
Figure 3: Multi-needle retrieval accuracy averaged over different needle insertion depths. Pretrained models are evaluated with 2K context (left), while finetuned models are evaluated with 4K context (right).
@@ 4.5 Kernel patterns
Other kernels, for example, can be viewed as priming, amplifying if the same key was attended by previous queries (head 5 on Figure 10), or edge detecting, amplifying the first or last of multiple contiguous keys with high attention (heads 8 and 11 on Figure 12). We leave further exploration of key-query kernel maps and their meaning for future research.
@@ 4.6 Ablations
Head convolution
An ablation study on the head kernel size shows that increasing kernel size improves average validation perplexity, as shown in Figure 6 (left), allowing for better communication across heads.
Figure 6: Ablation on the head kernel size in MTA convolution (left), and scaling laws (right). We report average validation perplexity on SlimPajama.
@@
Key-query conv Head conv
Pre-sm Post-sm Pre-sm Post-sm Group norm PPL $\downarrow$
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ scalar gating 10.90
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ depth scaling 10.92
$\checkmark$ $\times$ $\checkmark$ $\checkmark$ scalar gating 10.95
$\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ 10.99
$\checkmark$ $\times$ $\checkmark$ $\checkmark$ depth scaling 10.99
$\checkmark$ $\times$ $\checkmark$ $\checkmark$ no scaling 11.03
$\checkmark$ $\times$ $\times$ $\checkmark$ depth scaling 11.09
$\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ 11.16
$\checkmark$ $\times$ $\times$ $\checkmark$ no scaling 11.13
$\checkmark$ $\times$ $\times$ $\checkmark$ layer-norm scaling 11.41
$c_{q}=4$, $c_{k}=9$$\times$ $\times$ $\times$ depth scaling 11.23
$c_{q}=6$, $c_{k}=11$$\times$ $\times$ $\times$ depth scaling 11.23
$c_{q}=8$, $c_{k}=13$$\times$ $\times$ $\times$ depth scaling 11.31
$\times$ $\checkmark$ $\times$ $\checkmark$ depth scaling 11.11
$\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ 11.19
$\checkmark$ $\times$ $\checkmark$ $\times$ depth scaling 11.10
$\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ 11.20
@@
Scaling laws
We experiment with models of various sizes: 300M, 550M, and 1B, with hyperparameters listed in Table 7. As shown in Figure 6 (right), we observe a consistent pattern across models of different sizes, where MTA provides superior performance.
@@ 5 Related work
There has been a number of attempts to focus the soft attention mechanism on important context. For example, modifications have been proposed to make softmax sharper. Martins & Astudillo (2016) propose to replace softmax with sparsemax for sparse activations. Veličković et al. (2024); Nakanishi (2025) propose Adaptive temperature and Scalable-Softmax that adjust exponential base in softmax attention. Irrelevant tokens can be altogether removed from memory (Sukhbaatar et al., 2021) to make attention easier. In Golovneva et al. (2024); Desrochers et al. (2024) the authors incorporate contextual information into position encodings to allow positions to be amplified by context. More recently, several noise canceling mechanisms have been proposed for attention. Talking heads attention (Shazeer et al., 2020) adds linear projections across the attention-heads dimension before and after the softmax operation. Diff transformer (Ye et al., 2024) uses differential amplifiers to focus attention to the relevant context, which is related to our head mixing step. Cang et al. (2025) further extends it by introducing normalization steps. OpAmp adaptation (Wu et al., 2025) propose adapters, which are fine-tuned with noisy context to enhance an LLMs’ denoising capability. Xu et al. (2024) modified attention so that keys and values can shift by one time step. This can be viewed as a special case of MTA where the convolution performs shifting in the key dimension.
@@
Convolution layers have been used with attention (Gehring et al., 2017; Wu et al., 2019), especially in vision Transformers, to decompose each image into a sequence of tokens with fixed length, and then apply multiple standard Transformer layers (Xiao et al., 2021; Wu et al., 2021). There have been some attempts to include convolution in language modeling as well. For example, Liu et al. (2018) use convolution to compress the keys and values in the multi-headed attention by a factor of 3 thus allowing to process sequences 3x in length. Liu & Lapata (2019) later augment this architecture with the ability to encode multiple documents in a hierarchical manner, and further improve on long-context summarization tasks. Later Subramanian et al. (2020) propose three hierarchical models to learn different levels of abstraction in language by interchanging casual transformer layers with convolutions or pooling, or compressing representations for tokens further away in the past. Gulati et al. (2020) proposed a Conformer architecture for speech recognition, where a convolution module is applied to multi-head self-attention output. This architecture was later used to encode speech in the Llama-3 herd of models (Grattafiori et al., 2024). (Zheng et al., 2024) apply a convolution on the attention weights in the key dimension that enhances Transformer’s extrapolation capabilities to long context. However, none of these methods apply convolutions across multiple queries, keys and heads to the attention weights.
@@ Appendix C Toy task details
Figure 7: Accuracy (%) on QA1-5 tasks in BabiLong benchmark. Note the random performance on QA3 is 16.67%, thus all models perform poorly on QA3. On other tasks, MTA demonstrates strong performance compared to baselines especially when there is lengthy distraction text (4K).
@@ Appendix D Language model training details
Training setup follows LLaMa architecture (Touvron et al., 2023), and includes RMSNorm pre-normalization (Zhang & Sennrich, 2019), SwiGLU activation function (Shazeer, 2020), and Rotary Embeddings (Su et al., 2024). Table 7 shows a detailed hyperparameter breakdown for Transformer model pretraining. All models were trained in the same setup, except Diff Transformer, which had twice less heads to account for doubled dimension.
@@
Parameter 300M model 550M model 880M model 1B model
Dimension, $D$ $1024$ $1280$ $1536$ $2048$
Layers $20$ $24$ $24$ $24$
Heads, $M$ $16$ $10$ $16$ $16$
RoPE theta $100,000$ $100,000$ $100,000$ $100,000$
Batch size (tokens) $262,144$ $262,144$ $262,144$ $524,288$
Context length $2048$ $2048$ $2048$ $2048$
Learning rate $41.5e-$ 4 $41.5e-$ 4 $41.5e-$ 4 $41.5e-$ 4
Weight decay $0.05$ $0.05$ $0.05$ $0.05$
Warm up steps $375$ $375$ $375$ $187$
Table 7: Hyperparameters for Transformer training.
@@ Appendix G Multi-needle evaluation results
Head kernel patterns for the model with $c_{h}$=2 are shown in Figure 15. They are simpler due to their small size. Besides an identity with scaling, a common pattern is contrasting: subtracting one attention weight from another. We also observe that the kernel scales increase with layers when there is no group normalization (see Figure 17). This is probably to compete with the residual stream, which gets larger with the model’s depth. However, this pattern is not present with group normalization because it will undo the effect of such scaling.
@@
Appendix I Finetuning with MTA
One natural question readers might ask is if MTA could be integrated into models that were already trained with standard attention? Would this require complete retraining, or could it be added through some form of adaptation? Architectually, MTA can be added as additional layers to already trained models, and weights can be updated with continual training. Since the main component of MTA is a convolution, we can initialize it to identity and insert it in an existing a Transformer layer. Such a modification will not change the output of the layer, so when we start training the model, it should maintain its performance. However, as the added convolution starts deviating from the identity state, it will allow the transformer to condition its attention on multiple keys and queries.
To understand how well pre-trained models can learn we perform preliminary experiments by finetuning our 1.4B model, as well as opensourced Llama models (Grattafiori et al., 2024). In all these experiments we used the same finetuning setup similar to Section 4.3, but kept the context length at 2048 tokens.
Validation perplexity results are reported in Table 10. We observe that all models finetuned with MTA were not only able to incorporate new kernels, but outperform baselines in terms of perplexity.
Pretraining Cont training arxiv book c4 cc github se wiki Avg PPL $\downarrow$
1.4B models
Transformer Transformer 4.54 12.87 19.21 13.77 3.99 9.49 10.96 10.69
Transformer MTA 4.51 12.78 19.09 13.67 3.95 9.41 10.87 10.61
MTA MTA 4.47 12.57 18.82 13.47 3.89 9.25 10.65 10.44
Llama 3 herd
Llama 3.2 1B Llama 3.2 1B 4.54 14.60 18.14 13.40 3.92 8.84 10.25 10.53
Llama 3.2 1B MTA 4.52 14.49 18.04 13.32 3.89 8.79 10.19 10.46
Llama 3.2 3B Llama 3.2 3B 4.12 12.08 15.58 11.50 3.37 7.52 8.35 8.93
Llama 3.2 3B MTA 4.11 12.03 15.51 11.46 3.35 7.48 8.31 8.89
Llama 3.1 8B Llama 3.1 B 4.00 11.04 15.04 10.99 3.29 7.33 8.00 8.53
Llama 3.1 8B MTA 3.98 10.97 14.97 10.94 3.27 7.29 7.97 8.48
Table 10: Validation perplexity on SlimPajama dataset after continuous training for 10.5B tokens on our 1.4B models, and continuous training for 5.3B tokens for Llama 3 models.